OpenStudy (anonymous):

DESPERATELY NEED PRECAL/TRIG HELP I need help with parametric conics how to convert from parametric to rectangular form - y = 1/8 (x-3)^2 + 5 x = 2 (y + 7)^2 -8

OpenStudy (anonymous):

relating to cos and sin

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

these look like they are already in rectangular form

OpenStudy (anonymous):

Well then I guess convert it the other way? So it's like _ cos (t) _

OpenStudy (anonymous):

and sin (t)

OpenStudy (anonymous):

THIS IS JUST A EXAMPLE ON HOW TO DO IT Graphic Representation of the Equations : 3y + 4x = 29 -3y + 2x = 1 Solve by Substitution : // Solve equation [2] for the variable x [2] 2x = 3y + 1 [2] x = 3y/2 + 1/2 // Plug this in for variable x in equation [1] [1] 4•(3y/2+1/2) + 3y = 29 [1] 9y = 27 // Solve equation [1] for the variable y [1] 9y = 27 [1] y = 3 // By now we know this much : x = 3y/2+1/2 y = 3 // Use the y value to solve for x x = (3/2)(3)+1/2 = 5 Solution : {x,y} = {5,3} Processing ends successfully

OpenStudy (misty1212):

ok lets look it up

OpenStudy (misty1212):

\[x = 2 (y + 7)^2 -8\] is a parabola

OpenStudy (misty1212):

opens to the right

OpenStudy (misty1212):

the way to write this conic section in parametric form is \[r=\frac{ed}{1-\cos(\theta)}\]

OpenStudy (misty1212):

\(e\) is the eccentricity, which for a parabola is \(1\)

OpenStudy (misty1212):

and \(d\) is the directrix

OpenStudy (anonymous):

THIS IS JUST A EXAMPLE ON HOW TO DO IT Graphic Representation of the Equations : 3y + 4x = 29 -3y + 2x = 1 Solve by Substitution : // Solve equation [2] for the variable x [2] 2x = 3y + 1 [2] x = 3y/2 + 1/2 // Plug this in for variable x in equation [1] [1] 4•(3y/2+1/2) + 3y = 29 [1] 9y = 27 // Solve equation [1] for the variable y [1] 9y = 27 [1] y = 3 // By now we know this much : x = 3y/2+1/2 y = 3 // Use the y value to solve for x x = (3/2)(3)+1/2 = 5 Solution : {x,y} = {5,3} Processing ends successfully